# Understanding Sampler Stochasticity in Training Diffusion Models for RLHF

### v1

@@ 1 Introduction
Diffusion models (e.g., Stable Diffusion (Rombach et al., 2022), SDXL (Podell et al., 2024), FLUX (Black Forest Labs, 2024)) have shown strong performance in text-to-image (T2I) tasks, and have also been extended beyond images to video (Ho et al., 2022) and audio (Liu et al., 2023). To meet downstream objectives such as aesthetics, safety, and alignment, it is essential to post-train with RLHF (Ouyang et al., 2022) for preference-driven improvements, often with a KL-regularization term to preserve performance on pretrained tasks (Schulman et al., 2017). Widely used RLHF algorithms include DDPO (Black et al., 2024) and GRPO (Shao et al., 2024) variants (FlowGRPO (Liu et al., 2025), DanceGRPO (Xue et al., 2025), MixGRPO (Li et al., 2025)). DDPO directly optimizes human-preference rewards by casting the denoising process as a Markov Decision Process (MDP); GRPO variants use group-relative advantages. See (Winata et al., 2025) for a broader review of successful RLHF algorithms for generative models.
Despite strong progress in alignment, RLHF training often exhibits unstable trajectories, long inference times, and vulnerability to reward hacking (Skalse et al., 2022). Using multiple rewards can mitigate the latter (Lee et al., 2025), but we also need efficient, robust samplers to produce stable, high-quality fine-tuned models. A classical scheme is DDPM (Ho et al., 2020), a discretization of the score-based backward SDE (Risken, 1996; Song et al., 2021b), which preserves data marginals but is time-consuming. In contrast, the deterministic DDIM sampler (Song et al., 2021a) follows the probability-flow ODE, enabling marginal-preserving sampling with fewer steps. DDPO typically uses DDIM with constant stochasticity to generate training samples; MixGRPO mixes SDE and ODE steps, varying stochasticity across the denoising horizon. However, although stochasticity is valuable for generating diverse training data, fine-tuned models often use DDIM or higher-order ODE solvers (Lu et al., 2022; 2025) for fast, stable inference. This naturally raises the question:
Why can we guarantee good sampling quality when inference uses a different noise level than the one used during RLHF training?
SDE Refer to caption Refer to caption Refer to caption Refer to caption
ODE Refer to caption Refer to caption Refer to caption Refer to caption
In this paper, we address it from three perspectives. Theoretically, we derive non-vacuous bounds on the reward gap between a generally SDE-fine-tuned model and its ODE-sampling counterpart using Gronwall’s inequality. Specifically, for VE and VP Gaussian models, the gap shrinks at sharp rates $O(1/T)$ and $O(e^{-T^{2}}/2)$, where $T$ is the denoising time horizon. These results justify the practice of ODE inference for SDE-fine-tuned models and provide insights into more randomized exploration beyond $\eta=1.0$, the noise level for continuous score-based backward dynamics. Methodologically, we adopt the gDDIM (Zhang et al., 2023) framework for arbitrary stochasticity levels. Rather than linear variance interpolations between DDIM and DDPM, our implementation preserves marginals for all $\eta\geq 0$ and, more importantly, supports $\eta>1$ in a principled way, enabling broader RLHF exploration beyond standard DDPM noise levels. Empirically, we evaluate large-scale T2I models and RLHF algorithms to quantify the reward gap across multiple reward functions. The results show that (i) reward gaps decrease as training quality improves, (ii) moderate-to-high training stochasticity (e.g., $\eta=1.2$) yields superior in- and out-of-domain performance, and (iii) ODE inference remains stable and often outperforms SDE inference under small denoising step budgets.
  • Theory: Sharp bounds for VE/VP models; non-vacuous bounds for general distributions (linear in stochasticity, bounded in time-steps).
  • Methodology: gDDIM-based noise scheduling that preserves marginals for all noise levels ($\eta\geq 0$) and supports high stochasticity ($\eta>1$) in a principled way.
  • Experiments: Comprehensive evaluations across rewards and RLHF algorithms, validating bounded reward gaps and improved ODE inference with higher training stochasticity.
@@
Related Literature. Higher-order ODE solvers such as RX-DPM (Choi et al., 2025), DEIS (Zhang & Chen, 2022) approximate probability-flow. Recent analyses give broader convergence guarantees (Huang et al., 2025) and better representability (Chen et al., 2023) for score-based diffusion. In parallel, RL-based fine-tuning leverages stochastic sampling: DRaFT differentiates through noisy trajectories (Clark et al., 2023), Score-as-Action casts fine-tuning as stochastic control (Zhao et al., 2025), and Adjoint Matching enforces optimal memoryless noise schedules (Domingo-Enrich et al., 2024). Large-scale studies combine multiple rewards (Zhang et al., 2024); SEPO, D3PO, and ImageReFL boosts alignment (Zekri & Boullé, 2025; Yang et al., 2024; Sorokin et al., 2025). Finally, (Liang et al., 2025) gives discretization error bound and (Wu et al., 2024) discusses guidance (Ho & Salimans, 2021) for Gaussian Mixture priors.
Organization of the paper. The remainder of the paper is organized as follows. In Section 2, we provide background on diffusion models and the RLHF framework. The theory is developed in Sections 3 and 4. In Section 5, we conduct numerical experiments to justify the role of sampler stochasticity in fine-tuning large-scale T2I models. Concluding remarks are in Section 6.
@@ 2 Preliminaries
Figure 2: One-dimensional VP dynamics under different noise levels ($\eta$), using the same control function. Increasing $\eta$ introduces higher stochasticity in the trajectories while preserving the underlying dynamics.
@@ 2.1 Forward-Backward Diffusion Models
The continuous-time diffusion models (Song et al., 2021b) involve forward and backward procedures, and provide a framework that both unifies and extends earlier discrete-time models. The brief review below of this approach is based on (Tang & Zhao, 2025).
The goal of diffusion models is to generate new samples that resemble the target data distribution $p_{\mbox{data}}(\cdot)$. The forward process is given by an SDE:
@@
where $\{B_{t}\}$ is the $d$-dimensional Brownian motion, and $f:\mathbb{R}_{+}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ are given functions (model parameters). Some regularity conditions on $f(\cdot,\cdot)$ and $g(\cdot)$ are required to ensure equation 1 is well-defined, see (Stroock & Varadhan, 1979). Moreover, assume the forward process $\{X_{t}\}$ has a smooth density $p(t,x):=\mathbb{P}(X_{t}\in dx)/dx$. It is known (Anderson, 1982; Haussmann & Pardoux, 1986) that under time reversal, the process, denoted $\overline{X}_{t}=X_{T-t}$, is governed by the SDE:
$d\overline{X}_{t}=(-f(T-t,\overline{X}_{t})+g^{2}(T-t)\nabla\log p(T-t,\overline{X}_{t}))dt+g(T-t)dB_{t},\;\overline{X}_{0}\sim p(T,\cdot).$
There are two key points to note. First, as the goal is to generate the target distribution, which is unknown, the backward process is initiated with a prior $p_{\mbox{noise}}(\cdot)$, a proxy of $p(T,\cdot)$), instead of setting $\overline{X}_{0}\sim p(T,\cdot)$. The choice of $p_{\mbox{noise}}(\cdot)$ depends on each specific model, and varies case by case in applications. Similarly, the choice of the pair of functions $(f,g)$ also depends on applications, and notable examples include variance exploding (VE) and variance preserving (VP) SDEs (see Section 3).
Second, since the term $\nabla\log p(T-t,\overline{X}_{t})$ in the time-reversal process $\{\overline{X}_{t}\}$ is unknown, we need to find a way to estimate/approximate it. Specifically, calling $\nabla\log p(t,x)$ the “score function”, score matching refers to estimate the score function via a family of parameterized functions $\{s_{\theta}(t,x)\}_{\theta}$ (e.g., via neural nets). The most widely used approach is denoising score matching Vincent (2011):
$\min_{\theta}\mathbb{E}_{t\sim\mbox{Unif}\,[0,T],\,X_{0}\sim p_{data}(\cdot)}[\mathbb{E}_{p(t,\cdot|X_{0})}|s_{\theta}(t,X_{t})-\nabla\log p(t,X_{t}|X_{0})|^{2}].$
Replacing $\nabla\log p(t,x)$ by $s_{\theta}(t,x)$, the (backward) diffusion sampler is:
@@
For richer levels of trajectory variation while keeping terminal marginals, the sampler is set to:
@@
where $\eta\geq 0$ controls the sampler stochasticity along the trajectories. When $\eta=1$, equation 3 reduces to equation 2. When $\eta=0$, equation 3 becomes the probability flow ODE. We will use the processes specified by equation 3 as the pretrained models for fine-tuning. In particular, while prior studies focus on $\eta\in[0,1]$, we will explore the possible usage of $\eta>1$.
@@ 2.2 Discretization – Stochastic gDDIM
As shown in (Song et al., 2021b), the DDPM scheme is a discretized SDE that preserves variance, and the deterministic DDIM is a discretization of the probability flow ODE that preserves variance. For most diffusion models, the model parameters are:
@@
where $\{\alpha_{t}\}_{t=0}^{T}$ is a decreasing sequence with $\alpha_{0}=1$ and $\alpha_{T}=0$. (Zhang et al., 2023) proposed the following discretization of equation 3 for all $\eta\geq 0$:
@@
equation 4 is referred to as the generalized denoising diffusion implicit model (gDDIM), which agrees with the stochastic DDIM (Song et al., 2021a). The difference lies in the choice of $\sigma_{t}(\eta)$: (Song et al., 2021a) relied on heuristics to suggest $\sigma_{t}(\eta)\propto\eta$, whereas equation 5 proposed by (Zhang et al., 2023) guarantees the exact sampling if the score function is known/precise. When $\eta=0$, we have $\sigma_{t}(0)=0$, and equation 4 is just the deterministic DDIM. We use gDDIM scheme to conduct numerical experiments with controlled stochasticity, of which the results are detailed in Section 5.
@@ 2.3 Diffusion RLHF
RLHF was originally proposed for LLM alignment (Bai et al., 2022; Ouyang et al., 2022). Such a paradigm can also be made well-suited for fine-tuning diffusion models, particularly to enhance T2I generation aligned with human feedback. The first works in this direction are (Black et al., 2024; Fan et al., 2023; Lee et al., 2023), which formulated the denoising step as Markov decision processes (MDPs). (Gao et al., 2024; Zhao et al., 2024; 2025) developed an RL approach based on continuous-time models. Here we focus on the most widely used approach of Denoising Diffusion Policy Optimization (DDPO).
DDPO (Black et al., 2024): We formulate the denoising steps $\{x_{T},x_{T-1},\cdots,x_{0}\}$ as an MDP:
@@
where $p_{\theta}$ is the probability from $x_{t}$ to $x_{t-1}$ given prompt $\bm{c}$ and $\theta$-parametrized score function $s_{\theta}$; $\delta_{\bullet}$ is the Dirac function that distributes the (non-zero) mass at $\bullet$ only; $R$ is the reward function that is equal to the denoised image quality $r(x_{0},c)$ for each trajectory. The objective of DDPO is to maximize:
@@
The policy gradient of equation 6 is $\nabla_{\theta}\mathcal{J}_{DDPO}=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},\bm{c})\,r(x_{0},\bm{c})\right]$ (Williams, 1992). It is common to parameterize $p_{\theta}(x_{t-1}|x_{t},\bm{c})$ as isotopic Gaussian, so $\nabla_{\theta}\mathcal{J}_{DDPO}$ can be easily computed by Monte Carlo (Mohamed et al., 2020).
GRPO (Shao et al., 2024): A recent breakthrough in the space of RLHF is the Group Relative Policy Optimization (GRPO) framework in the wake of Deepseek-R1 for reasoning. For T2I tasks, GRPO uses a group of sampled outputs (indexed by $i$ or $k$ below), along with the group normalization/average per prompt, to compute the advantages:
@@
where $G$ is the group size; $\text{std}\left(\{r(x_{0}^{k},\bm{c})\}_{k=1}^{G}\right)$ denotes the standard deviation of the group rewards. The objective of GRPO is to maximize:
$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=1}^{T}\,\min\left(\rho_{t}^{i}(\theta)A_{i},\text{clip}(\rho_{t}^{i}(\theta),1-\epsilon,1+\epsilon)A_{i}\right)\right],$ (7)
where $\rho_{t}^{i}(\theta)=\frac{p_{\theta}(x_{t-1}^{i}|x_{t}^{i},\bm{c})}{p_{\theta_{\mbox{old}}}(x_{t-1}^{i}|x_{t}^{i},\bm{c})}$ is the likelihood ratio; $\{x_{t}^{i}\}$ is the $i$-th sample trajectory in the group; and the expectation is with respect to $\bm{c}\sim p(\bm{c})$ and $\{x^{i}_{t}\}\sim\pi_{\theta_{\mbox{old}}}(\cdot|\bm{c})$. Variants of the GRPO algorithm (Xue et al., 2025; Li et al., 2025) have been proposed for T2I tasks.
@@ 3 Theory on the reward gap
As mentioned in the last section, the forward-backward diffusion approach to generative AI usually makes use of stochasticity for exploration (the forward step), while relying on deterministic samplers for inference (the backward step); the latter is obviously motivated by computational efficiency. Yet, there’s no study to examine/analyze the role of stochasticity; namely, whether there’s a right or “best” level of stochasticity to use, in both steps; and in particular, whether it’d be worthwhile to use stochasticity to the backward step, at the price of reduced efficiency.
Denote by $\{Y^{REF}_{t}\}$ the “reference” model specified in equation 3), with a specific choice of the parameter $\eta$. Denote by $\{Y^{SDE}_{t}\}:=\{Y^{\theta}_{t}\}$ the fine-tuned model, score functions parameterized by $\theta$.
To carry out this analysis, consider the following entropy-regularized reward objective (e.g., (Uehara et al., 2024; Tang, 2024)):
$F_{\eta}(\theta)=\mathbb{E}[r(Y^{\theta}_{T})]-\beta\,\text{KL}(Y_{T}^{\theta}||Y_{T}^{REF}),$ (8)
where $r(\cdot)$ is a reward function, and $\beta>0$ controls the level of exploration relative to the pretrained model. $Y^{SDE}_{t}$ evolves under a same $\eta$ as $Y^{REF}_{t}$, and the problem is to solve $\theta^{*}_{\eta}:=\arg\max_{\theta}F_{\eta}(\theta)$.
$\bullet$ Inference: Denote by $\{Y_{t}^{ODE}\}$ the marginal-preserved (deterministic) ODE sampler by letting $\eta\downarrow 0$ in $\{Y^{SDE}_{t}\}$.
Also let $\mathcal{J}_{\bullet}:=\mathbb{E}[r(Y_{T}^{\bullet})]$ be the evaluation metric, with $\bullet\in\{REF,SDE,ODE\}$. The goal is to understand how much improvement gained from fine-tuning, and the reward gap induced by the (deterministic) ODE sampler. This motivates the following definitions:
@@
    Improvement of fine-tuning as the reward difference between $\{Y^{ODE}_{t}\}$ and $\{Y^{REF}_{t}\}$, $\mathcal{}I_{\eta}:=\mathcal{J}_{ODE}-\mathcal{J}_{REF}\geq 0$.
@@
    Reward gap as the reward difference between $\{Y^{ODE}_{t}\}$ and $\{Y^{SDE}_{t}\}$, $\Delta_{\eta}:=|\mathcal{J}_{ODE}-\mathcal{J}_{SDE}|$.
In what follows, we study $I_{\eta}$ and $\Delta_{\eta}$ for the VE/VP models with a Gaussian or mixture Gaussian target distribution, where the score function has a closed-form expression. A non-vacuous bound for the general case will be given in the next section.
@@
$\mathcal{J}_{ODE}\geq\mathcal{J}_{REF}$ for all models, following from the optimality of $\theta^{*}_{\eta}$. $\mathcal{J}_{ODE}>\mathcal{J}_{SDE}$ for VE/VP models and DDPO; $\mathcal{J}_{SDE}>\mathcal{J}_{ODE}$ for MixGRPO.

### v2

@@ 1 Introduction
Diffusion models (e.g., Stable Diffusion (stabledif), SDXL (sdxl), FLUX (flux)) have shown strong performance in text-to-image (T2I) tasks, and have also been extended beyond images to video (video) and audio (audio). To meet downstream objectives such as aesthetics, safety, and alignment, it is essential to post-train with RLHF (rlhf) for preference-driven improvements, often with a KL-regularization term to preserve performance on pretrained tasks (ppo). Widely used RLHF algorithms include DDPO (ddpo) and GRPO (grpo) variants (FlowGRPO (FlowGRPO), DanceGRPO (DanceGRPO), MixGRPO (MixGRPO)). DDPO directly optimizes human-preference rewards by casting the denoising process as a Markov Decision Process (MDP); GRPO variants use group-relative advantages. RLHF training may also exhibit unstable trajectories, long inference times, and vulnerability to reward hacking (rewardhack; CMPO) and efficient, robust samplers (lu2022dpm; lu2025dpm) are desired for stable, high-quality fine-tuned models. See (winata2025preference) for a broader review of successful RLHF algorithms for generative models.
The classical scheme DDPM (ddpm), a discretization of the score-based backward SDE (fp; Song20), allows for trajectory diversity and rich exploration in RLHF training; the deterministic DDIM sampler (ddim) follows a probability-flow ODE and facilitates inference. High stochasticity during training is crucial for effective exploration of the reward landscape. However, during inference, deterministic ODE sampling is preferred to ensure consistency and computational efficiency. This creates an inherent tension: we train via a highly stochastic process (SDE) but we deploy a deterministic process (ODE) for inference. This discrepancy necessitates the following question:
To what extent can we guarantee ODE inference quality after we fine-tune diffusion models with SDE sampled trajectories?
Figure 1: An illustration of the problem formulation. $\Delta_{\eta}$ represents the “Reward Gap” and $I_{\eta}$ represents the ODE reward improvement after SDE fine-tune. We would like $\Delta_{\eta}\downarrow$ with $I_{\eta}\uparrow$.
Figure 1 illustrates the problem formulation. While SDE and ODE denoising processes (equation 3) yield the same terminal marginal for the pretrained model (see Section 2.1), a distributional difference arises between the distribution ($p_{\text{data}}^{\text{SDE}}$) after stochastic fine-tuning and the deterministic inference outcome ($p_{\text{data}}^{\text{ODE}}$) (see Section 3). We denote $\eta$ as the stochasticity scale, where $\eta=1.0$ corresponds to the standard variance in DDPM. We define this discrepancy as the Reward Gap ($\Delta_{\eta}$), which must be bounded to preserve the ODE reward improvement ($I_{\eta}$) gained over the baseline. In this paper, we present our analysis from three perspectives:
  • Theory: We bound $\Delta_{\eta}$ between a generally SDE-fine-tuned model and its ODE-sampling counterpart using Gronwall’s inequality. Specifically, for VE, VP and mixture Gaussian models, the gap shrinks at sharp rates $O(1/T)$ and $O(e^{-T^{2}}/2)$, where $T$ is the denoising time horizon.
  • Methodology: To handle high stochasticity beyond $\eta=1.0$ without changing marginals, we adopt the gDDIM scheme (gddim) for arbitrary stochasticity levels. This framework is introduced in Section 2.2 and 5. See Appendix E for more technical details.
  • Experiments: We evaluate large-scale T2I models and RLHF algorithms, DDPO and MixGRPO, to quantify the reward gap across multiple reward functions. The results show that $\Delta_{\eta}$ decrease as training quality improves and $\eta=1.2$ yields superior ODE inference performance, $I_{\eta}$.
@@
Related Literature. Higher-order ODE solvers such as RX-DPM (choi2025enhanced), DEIS (zhang2022fast) approximate probability-flow. Recent analyses give broader convergence guarantees (huang2025fast) and better representability (chen2023sampling) for score-based diffusion. In parallel, RL-based fine-tuning leverages stochastic sampling: DRaFT differentiates through noisy trajectories (clark2023directly), Score-as-Action casts fine-tuning as stochastic control (ZC25), and Adjoint Matching enforces optimal memoryless noise schedules (domingo2024adjoint). Large-scale studies combine multiple rewards (zhang2024large); SEPO, D3PO, and ImageReFL boosts alignment (zekri2025fine; yang2024using; sorokin2025imagerefl). Finally, (gaussianmixture) gives discretization error bound and (gm) discusses guidance (cfg) for Gaussian Mixture priors.
Organization of the paper. The remainder of the paper is organized as follows. In Section 2, we provide background on diffusion RLHF and gDDIM in particular. Section 3 computes exact $\Delta_{\eta}$ and $I_{\eta}$ for simple controls. More complex networks are analyzed in Section 4. Numerical experiments in Section 5 validate our theoretical bounds by fine-tuning large-scale T2I models with DDPO and GRPO algorithms. Conclusion is given in Section 6.
@@ 2.1 Forward-Backward Diffusion Models
The goal of diffusion models (Song20) is to generate new samples similar to $p_{\mbox{data}}(\cdot)$. The forward process is given by an SDE:
@@
where $\{B_{t}\}$ is the $d$-dimensional Brownian motion, and $f:\mathbb{R}_{+}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ are given model parameters under regularity conditions (SV79).
We use denoising score matching (Vi11) to minimize the mean square error between a parametrized neural network $s_{\theta}(t,X_{t})$ and the unknown term $\nabla\log p(T-t,\overline{X}_{t})$ in the time reversal SDE (HP86). The reverse-time SDE becomes:
@@
For richer levels of sampler stochasticity, a flexible stochasticity parameter $\eta\geq 0$ is introduced:
@@
With a divergence-free noise scale, equation 3 preserves the terminal marginals (Ander82).
@@ 2.2 Discretization – Stochastic gDDIM
As shown in (Song20), for most diffusion models, the model parameters are:
@@
where $\{\alpha_{t}\}_{t=0}^{T}$ is a decreasing sequence from 1 to 0. For all $\eta\geq 0$, (gddim) proposed the generalized DDIM update:
@@
For any $\eta>0$, the discrete scheme equation 4 has terminal marginal equal to equation 3. We use this discretization formulation for post-training with $\eta>1.0$ in Section 5. Details in Appendix E.
@@ 2.3 Diffusion RLHF
RLHF was originally proposed for LLM alignment (Bai22; rlhf). (ddpo; dpok; Lee23) first formulated the denoising step as Markov decision processes (MDPs) to fine-tune diffusions. (Gao24; ZC24; ZC25) developed an RL approach based on continuous-time models. Here we focus on Denoising Diffusion Policy Optimization (DDPO) and Group Relative Policy Optimization (GRPO).
DDPO (ddpo): We formulate the denoising steps $\{x_{T},x_{T-1},\cdots,x_{0}\}$ as an MDP:
@@
where $\delta_{\bullet}$ is the Dirac point mass concentrating at $\bullet$. The objective of DDPO is to maximize:
@@
The policy gradient of equation 6 is $\nabla_{\theta}\mathcal{J}_{DDPO}=\mathbb{E}\left[\sum_{t=0}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},\bm{c})\,r(x_{0},\bm{c})\right]$ (reinforce). A common parameterization of $p_{\theta}$ is isotropic Gaussian (montecarlo).
GRPO (grpo): GRPO first computes the advantages:
@@
where $G$ is the group size; $\text{std}\left(\{r(x_{0}^{k},\bm{c})\}_{k=1}^{G}\right)$ denotes the standard deviation of group rewards. The objective of GRPO is to maximize:
$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{N}\sum_{n=1}^{N}\,\min\left(\rho_{t}^{i}(\theta)A_{i},\text{clip}(\rho_{t}^{i}(\theta),1-\epsilon,1+\epsilon)A_{i}\right)\right],$ (7)
where $\operatorname{clip}(\rho_{t}^{i}(\theta),1-\epsilon,1+\epsilon):=\min\{\max\{\rho_{t}^{i}(\theta),1-\epsilon\},\,1+\epsilon\}$ restricts the likelihood ratio $\rho_{t}^{i}(\theta)=\frac{p_{\theta}(x_{t-1}^{i}|x_{t}^{i},\bm{c})}{p_{\theta_{\mbox{old}}}(x_{t-1}^{i}|x_{t}^{i},\bm{c})}$ within the range $1-\epsilon,1+\epsilon]$ by imposing hard constraints on the boundaries; $\{x_{t}^{i}\}$ is the $i$-th sample trajectory in the group; $N$ is the number of grouped outputs, and the expectation is with respect to $\bm{c}\sim p(\bm{c})$ and $\{x^{i}_{t}\}\sim\pi_{\theta_{\mbox{old}}}(\cdot|\bm{c})$.
@@ 3 Theory on the reward gap
As discussed in Section 1, forward–backward diffusion models typically employ stochastic SDE dynamics for exploration in the RLHF training, while adopting deterministic ODE samplers for inference (see Figure 1). In practice, we follow entropy-regularized objectives to fine-tune a parametrized family of score functions (UZ24; Tang24):
@@
    $\{Y_{T}^{\mathrm{REF}}\}$ as the samples extracted from the referenced base model following equation 3 with terminal marginal $p_{data}$;
@@
    We fine-tune the following objectives under a fixed stochasticity parameter $\eta$:
    $F_{\eta}(\theta)=\mathbb{E}[r(Y_{T}^{\theta})]-\beta\,\mathrm{KL}(Y_{T}^{\theta}\,\|\,Y_{T}^{\mathrm{REF}}),\qquad\theta_{\eta}^{*}:=\arg\max_{\theta}F_{\eta}(\theta);$ (8)
    let $\{Y_{t}^{\mathrm{SDE}}\}=\{Y_{t}(\theta_{\eta}^{*})\}$ be the optimal fine-tuned model, and $\{Y_{t}^{\mathrm{ODE}}\}$ be the deterministic sampler obtained by letting $\eta=0$:
    $dY_{t}^{\text{SDE}}=\left(-f(T-t,Y_{t}^{\text{SDE}})+\frac{1+\eta^{2}}{2}g^{2}(T-t)s_{\theta_{\eta}^{*}}(T-t,Y_{t}^{\text{SDE}})\right)dt+\eta\,g(T-t)dB_{t},$ (9)
    $dY_{t}^{\text{ODE}}=\left(-f(T-t,Y_{t}^{\text{ODE}})+\frac{1}{2}g^{2}(T-t)s_{\theta_{\eta}^{*}}(T-t,Y_{t}^{\text{ODE}})\right)dt.$
Although a pretrained model admits identical terminal marginals for arbitrary $\eta$ in the form of equation 3, once the data distribution is aligned to a human-preference reward $r(x)$, the distribution of $\{Y_{T}^{\mathrm{ODE}}\}$ and $\{Y_{T}^{\mathrm{SDE}}\}$ (we denote as $p_{data}^{\text{SDE}}$ and $p_{data}^{\text{ODE}}$) are not necessarily the same.
Under a fixed noise level $\eta$, the terminal marginal satisfies a reward-tilting distribution $p_{data}^{\text{SDE}}\propto p_{\text{data}}(x)\exp(r(x)/\beta)$ (zhang2024rewardoveropt). The terminal score function
$s_{\theta_{\eta}^{*}}(T,x)\approx\nabla\log p_{\theta_{\eta}^{*}}(T,x)=\nabla\log p_{\text{data}}(x)+\frac{\nabla r(x)}{\beta},$
but the fine-tuned parameter $\theta_{\eta}^{*}$ does not regulate the intermediate trajectory, as the RLHF objective equation 8 only considers the terminal marginal. Therefore, we could not establish a general relationship between $s_{\theta_{\eta}^{*}}(t,x)$, $r(x)$, and $\nabla\log p_{\theta_{\eta}^{*}}(t,x)$ for $0<t<T$. In general,
$\nabla\!\cdot\left(p_{\theta_{\eta}^{*}}(t,x)\cdot s_{\theta_{\eta}^{*}}(t,x)\right)\neq\Delta p_{\theta_{\eta}^{*}}(t,x).$
As the reverse-time Fokker–Planck correspondence no longer satisfies the divergence free condition (Ander82), $p_{data}^{\text{SDE}}$ and $p_{data}^{\text{ODE}}$ need not coincide.
Figure 2: One-dimensional VP dynamics under different noise levels $\eta$, using the same control function, demonstrating $\mathbb{E}(Y_{T}^{\text{SDE}})\approx\mathbb{E}(Y_{T}^{\text{ODE}})$ despite $p_{data}^{\text{SDE}}\neq p_{data}^{\text{ODE}}$.
However, we are motivated by an observation in Figure 2 that 1D Variance Preserving (VP) dynamics, when run with the same constant downward drift, maintain a similar mean value regardless of the noise level $\eta$. we next rigorously quantify and generalize this observation Bounded reward gaps for simple dynamics are important for two reasons:
  • Error Isolation. Under VP (as well as VE) models, we are able to compute the exact reward difference induced by changing $\eta$, free from the well-studied score-approximation and discretization error (gaussianmixture), with which we will also analyze in Section 4 and upper-bound reward gaps in those more complex settings.
  • Parametrization Insight. Parametrization is inherent to the pre-trained (see equation 2) and the fine-tuned model (see equation 9) and affect RLHF performances in practice (han2025stochastic). Here we show that reward gap is small even under simple parametrization. As Section 4 bounds complex controls quadratic in $\eta$, our results are valid for general networks.
Our goals are to quantify the reward improvement of fine-tuning $I_{\eta}$ and the reward gap $\Delta_{\eta}$ between SDE and ODE inference. Here we formally define:
Definition 3.2.
As illustrated in Figure 1, we define
  1. 1.
    Improvement of fine-tuning as the reward difference between $\{Y^{\text{ODE}}_{t}\}$ and $\{Y^{\text{REF}}_{t}\}$, $\mathcal{}I_{\eta}:=\mathbb{E}[r(Y_{T}^{\text{ODE}})]-\mathbb{E}[r(Y_{T}^{\text{REF}})]\geq 0.$
  2. 2.
    Reward gap as the reward difference between $\{Y^{\text{ODE}}_{t}\}$ and $\{Y^{\text{SDE}}_{t}\}$, $\Delta_{\eta}:=\left|\mathbb{E}[r(Y_{T}^{\text{ODE}})]-\mathbb{E}[r(Y_{T}^{\text{SDE}})]\right|.$
In what follows, we analyze VE/VP models with a Gaussian or mixture Gaussian target distribution, where the score function has a closed-form expression without approximation and discretization errors, so the only discrepancy comes from changing $\eta$.
@@
$\mathbb{E}[r(Y_{T}^{\text{ODE}})]\geq\mathbb{E}[r(Y_{T}^{\text{REF}})]$ holds for all models, following from the optimality of $\theta^{*}_{\eta}$. However, the sign of $\left(\mathbb{E}[r(Y_{T}^{\text{ODE}})]-\mathbb{E}[r(Y_{T}^{\text{SDE}})]\right)$ depends on the process dynamics as well as RLHF algorithms.